# Implicit and Explicit Language Guidance for Diffusion-based Visual Perception

### v1

@@ Abstract
Text-to-image diffusion models have shown powerful ability on conditional image synthesis. With large-scale vision-language pre-training, diffusion models are able to generate high-quality images with rich texture and reasonable structure under different text prompts. However, it is an open problem to adapt the pre-trained diffusion model for visual perception. In this paper, we propose an implicit and explicit language guidance framework for diffusion-based perception, named IEDP. Our IEDP comprises of an implicit language guidance branch and an explicit language guidance branch. The implicit branch employs frozen CLIP image encoder to directly generate implicit text embeddings that are fed to diffusion model, without using explicit text prompts. The explicit branch utilizes the ground-truth labels of corresponding images as text prompts to condition feature extraction of diffusion model. During training, we jointly train diffusion model by sharing the model weights of these two branches. As a result, implicit and explicit branches can jointly guide feature learning. During inference, we only employ implicit branch for final prediction, which does not require any ground-truth labels. Experiments are performed on two typical perception tasks, including semantic segmentation and depth estimation. Our IEDP achieves promising performance on both tasks. For semantic segmentation, our IEDP has the mIoU score of 55.9% on AD20K validation set, which outperforms the baseline method VPD by 2.2%. For depth estimation, our IEDP outperforms the baseline method VPD with a relative gain of 10.2%.
@@
Diffusion models have exhibited exceptional performance in the related field of image synthesis [1, 2, 3], which can recover high-quality images from noisy data. To reduce the significant training cost in pixel space, stable diffusion model [4] has been proposed to perform diffusion forward and reverse backward on feature latent space. Through training on large-scale image-text paired dataset LAION-5B[5], stable diffusion model can generate various high-quality images with rich texture and reasonable structure under various text prompts. This phenomenon suggests that stable diffusion model has the strong feature representation ability of both low-level and high-level semantic information.
@@
Recently, the researchers [6, 7, 10] started to explore diffusion model for visual perception via fully utilizing its strong feature representation ability. VPD [6] first generates text prompts with all semantic classes existing in the dataset, and then utilizes these text prompts to condition feature extraction of stable diffusion. VPD has exhibited promising performance on various perception tasks, especially when using fewer training iterations. However, VPD employs all semantic classes from the entire dataset, which does not align the specific image that only contains several classes. To address this issue, TADP [7] adopts the pre-trained vision-language model BLIP-2 [8] to generate image-aligned captions, and employs these image-aligned captions to condition feature extraction of diffusion model. Compared to VPD, TADP avoids the interference of non-existing classes when conditioning feature extraction in some degree. We summarize the pipelines of VPD and TADP into Fig. 1(a). They initially rely on manual design or automatic generation to create text prompts. Subsequently, they utilize a CLIP text encoder to extract text embeddings, and fed them into stable diffusion to condition feature extraction. We argue that there exists some limitations in these methods. First, during inference, it is relatively cumbersome in TADP to employ two models to generate text prompts and text embeddings, and there exists a misalignment issue between text prompt and image in VPD. Second, during training, these methods do not fully utilize ground-truth labels corresponding to specific training images, which accurately align with the training images.
@@
To address these issues above, we propose an implicit and explicit language guidance framework for diffusion-based perception, named IEDP. Our proposed IEDP contains two different branches: implicit language guidance branch and explicit language guidance branch. In the implicit branch, we introduce a novel implicit prompt module that directly generates implicit text embeddings using CLIP image encoder, instead of using explicit text prompts. The implicit text embeddings are fed to stable diffusion to condition feature extraction. Because CLIP model is able to learn connection between image-text pairs, the implicit text embeddings contains class information existing in the image. In the explicit branch, we introduce an explicit prompt module to directly employ ground-truth labels of corresponding training image as explicit text prompts. The ground-truth labels provide accurate class information of the image, which can better guide feature learning. As shown in Fig. 1(b), during training, we jointly train the network via weight sharing between these two branches. During inference, we only employ implicit branch for perception prediction, and remove the explicit branch since the accurate ground-truth labels of test image are not available. We perform the experiments on typical visual perception tasks, including semantic segmentation and depth estimation. The experimental results demonstrate the efficacy of our proposed method. The contributions and merits of our proposed method can be summarized as follows.
@@
Our method achieves superior performance on semantic segmentation and depth estimation. For example, our IEDP achieves mIoU score of 55.9% on ADE20K[11] set, outperforming diffusion-based method VPD [6] by 2.2%. On depth estimation, our IEDP achieves RMSE of 0.228 on NYUv2[12], outperforming VPD by 10.2% relatively.
@@ II-A Visual Perception
As a typical perception task, semantic segmentation aims to group the pixels in an image into different semantic categories. In past years, semantic segmentation has achieved remarkable progress, which can be mainly divided into CNN-based and transformer-based approaches. Most CNN-based approaches were developed on fully convolutional framework [13]. To improve segmentation performance, some researchers explored to exploit local contextual information using encoder-decoder [14, 15, 16] or spatial pyramid structure [17, 18], while some other researchers proposed to exploit non-local contextual information using attention mechanism [19, 20, 21]. Recently, with the success of transformer in image classification [22], the researchers started to explore transformer-based approaches. Some approaches replace the convolutional backbone with the transformer backbone [23, 24, 25], while some other approaches design query-based transformer decoder for semantic segmentation [26, 27].
@@ IV-A Datasets and Evaluation Metrics
NYUv2. The NYUv2 dataset [12] is a typical depth estimation dataset collected from a variety of indoor scenes. The images and corresponding depth maps are captured by the pairs of RGB and Depth cameras. There are 24 images for training, and 645 images for testing. We adopt three commonly used indicators for depth estimation evaluation, including absolute relative error (REL), root mean squared error (RMSE), and average error between predicted depth and ground-truth depth . We further present threshold accuracy which represents =% of pixels satisfying for .
@@ IV-B Experiment Setups
Depth estimation. We adopt the similar decoder as in [47]. During training, we adopt AdamW as the optimizer, and employ the poly learning rate rate schedule. We randomly crop the images into 480480 pixels, and set the batch size as 24. There are totally 25 epochs The initial learning rate is 5e-4 and the weight decay as 0.1. Because depth estimation dataset does not provide ground-truth labels, we only employ implicit language guidance branch when performing depth estimation. During inference, we employ the techniques of horizontal flipping and sliding windows as in VPD.
@@
TADP[7] | ArXiv’23 | 0.225 | 0.976 | 0.997 | 0.999 | 0.062 | 0.027 |
IEDP (Ours) | - | 0.228 | 0.975 | 0.997 | 0.999 | 0.063 | 0.027 |
@@ IV-C State-of-the-art Comparison
Depth estimation. We compare our proposed method with some state-of-the-art methods on NYUv2 dataset in Table II. These state-of-the-art methods are also divided into classical approaches and diffusion-based approaches. Compared to diffusion-based approaches, our IEDP achieves promising performance. For example, DDP [55] has the RMSE score of 0.329, VPD [6] has the RMSE score of 0.254, and our IEDP has the RMSE score of 0.228. Therefore, compared to DDP and VPD, our IEDP has the relative improvements of 30.7% and 10.2% in terms of RMSE. Compared to TADP [7], our proposed IEDP has a comparable performance, but does not requires additional BLIP-2 for generating text prompts.
@@
IV-D Ablation Study
Effectiveness of two proposed modules. We first perform the ablation study to show the effectiveness of implicit and explicit prompt modules. Table III presents the results of progressively integrating these two modules into the baseline. The baseline adopts the unaligned text prompts as in VPD [6], which has the mIoU score of 48.1%. When replacing the unaligned text prompts with our implicit prompts, it has the mIoU score of 49.8%, outperforming the baseline by 1.7%. When further integrating our explicit prompt module, it has the mIoU score of 52.5%, which outperforms the baseline by 4.4% totally.
Different designs of the adapter. Table IV shows the results of different adapter designs in implicit branch, where the adapter aims to convert the visual features of CLIP image encoder to implicit text embeddings for denoising UNet. We give two different designs, including using an MLP layer and using learnable queries to convert visual features. We observe that using learnable queries has the better performance.
Number of learnable queries. Table V shows the impact of different numbers of learnable queries in implicit branch. We observe that there exists a certain degree of performance degradation when using small number of queries, but the decline is relatively limited. Moreover, compared to the baseline, using different queries all have the improvements. We employ the number of queries as 256 as our default setting.
Impact of position embeddings. Table VI shows the impact of using position embeddings in both implicit and explicit branch. When not using the position embeddings, it has 0.2% performance drop on mIoU. Therefore, we add position embeddings in both two branches.
Ground-truth labels vs. BLIP-based captions. In explicit language guidance branch, instead of using ground-truth labels, we can also employ BLIP[56] to automatically generate image captions as text prompts like TADP. Table VII compares ground-truth labels and BLIP-based captions. We observe that it has better performance using ground-truth labels that are less noisy.
@@ V Conclusion
In this paper, we propose an implicit and explicit language guidance framework for diffusion-based visual perception, named IEDP. Our IEDP introduces an implicit prompt module and an explicit prompt module into text-to-image diffusion model. In implicit prompt module, we employ a frozen CLIP image encoder to directly generate implicit text embeddings and fed these embeddings to diffusion model to condition feature extraction. In explicit prompt module, we utilize the ground-truth labels of training images as explicit text prompts and employ CLIP text encoder to generate text embeddings for diffusion model. The implicit prompt module and explicit prompt module jointly train the diffusion model, and only implicit prompt module is used during inference. Experiments are performed on semantic segmentation and depth estimation tasks to show the efficacy of proposed method.

### v2

@@ Abstract
Text-to-image diffusion models have shown powerful ability on conditional image synthesis. With large-scale vision-language pre-training, diffusion models are able to generate high-quality images with rich texture and reasonable structure under different text prompts. However, it is an open problem to adapt the pre-trained diffusion model for visual perception. In this paper, we propose an implicit and explicit language guidance framework for diffusion-based perception, named IEDP. Our IEDP comprises an implicit language guidance branch and an explicit language guidance branch. The implicit branch employs frozen CLIP image encoder to directly generate implicit text embeddings that are fed to diffusion model, without using explicit text prompts. The explicit branch utilizes the ground-truth labels of corresponding images as text prompts to condition feature extraction of diffusion model. During training, we jointly train diffusion model by sharing the model weights of these two branches. As a result, implicit and explicit branches can jointly guide feature learning. During inference, we only employ implicit branch for final prediction, which does not require any ground-truth labels. Experiments are performed on two typical perception tasks, including semantic segmentation and depth estimation. Our IEDP achieves promising performance on both tasks. For semantic segmentation, our IEDP has the mIoU score of 55.9% on AD20K validation set, which outperforms the baseline method VPD by 2.2%. For depth estimation, our IEDP outperforms the baseline method VPD with a relative gain of 11.0%.
@@
Diffusion models have exhibited exceptional performance in the related field of image synthesis [1, 2, 3, 4, 5], which can recover high-quality images from noisy data. To reduce the significant training cost in pixel space, stable diffusion model [6] has been proposed to perform diffusion forward and reverse backward on feature latent space. Through training on large-scale image-text paired dataset LAION-5B[7], stable diffusion model can generate various high-quality images with rich texture and reasonable structure under various text prompts. This phenomenon suggests that stable diffusion model has the strong feature representation ability of both low-level and high-level semantic information.
@@
Recently, the researchers [8, 9] started to explore diffusion model for visual perception via fully utilizing its strong feature representation ability. VPD [8] first generates text prompts with all semantic classes existing in the dataset, and then utilizes these text prompts to condition feature extraction of stable diffusion. VPD has exhibited promising performance on various perception tasks, especially when using fewer training iterations. However, VPD employs all semantic classes from the entire dataset, which does not align the specific image that only contains several classes. To address this issue, TADP [9] adopts the pre-trained vision-language model BLIP-2 [10] to generate image-aligned captions, and employs these image-aligned captions to condition feature extraction of diffusion model. Compared to VPD, TADP avoids the interference of non-existing classes when conditioning feature extraction in some degree. We summarize the pipelines of VPD and TADP into Fig. 1(a). They initially rely on manual design or automatic generation to create text prompts. Subsequently, they utilize a CLIP text encoder to extract text embeddings, and fed them into stable diffusion to condition feature extraction. We argue that there exists some limitations in these methods. First, during inference, it is relatively cumbersome in TADP to employ two models to generate text prompts and text embeddings, and there exists a misalignment issue between text prompt and image in VPD. Second, during training, these methods do not fully utilize ground-truth labels corresponding to specific training images, which accurately align with the training images.
@@
To address these issues above, we propose an implicit and explicit language guidance framework for diffusion-based perception, named IEDP. Our proposed IEDP contains two different branches: implicit language guidance branch and explicit language guidance branch. In the implicit branch, we introduce a novel implicit prompt module that directly generates implicit text embeddings using CLIP image encoder, instead of using explicit text prompts. The implicit text embeddings are fed to stable diffusion to condition feature extraction. Because CLIP model is able to learn connection between image-text pairs, the implicit text embeddings contains class information existing in the image. In the explicit branch, we introduce an explicit prompt module to directly employ ground-truth labels of corresponding training image as explicit text prompts. The ground-truth labels provide accurate class information of the image, which can better guide feature learning. As shown in Fig. 1(b), during training, we jointly train the network via weight sharing between these two branches. During inference, we only employ implicit branch for perception prediction, and remove the explicit branch since the accurate ground-truth labels of test image are not available. We perform the experiments on typical visual perception tasks, including semantic segmentation and depth estimation. The experimental results demonstrate the efficacy of our proposed method. The contributions and merits can be summarized as follows.
@@
Our method achieves superior performance on semantic segmentation and depth estimation. Our IEDP achieves mIoU score of 55.9% on ADE20K[12] set, outperforming diffusion-based method VPD [8] by 2.2%. On depth estimation, our IEDP achieves RMSE of 0.226 on NYUv2[13], outperforming VPD by 11.0% relatively. In addition, compared to VPD and TADP, our IEDP has the better trade-off between performance and inference time.
@@ II-A Visual Perception
As a typical perception task, semantic segmentation [14, 15] aims to group the pixels in an image into different semantic categories. In past years, semantic segmentation has achieved remarkable progress, which can be mainly divided into CNN-based and transformer-based approaches. Most CNN-based approaches were developed on fully convolutional framework [16]. To improve segmentation performance, some researchers explored to exploit local contextual information using encoder-decoder [17, 18, 19] or spatial pyramid structure [20, 21], while some other researchers proposed to exploit non-local contextual information using attention mechanism [22, 23, 24]. Recently, with the success of transformer in image classification [25], the researchers started to explore transformer-based approaches. Some approaches replace the convolutional backbone with the transformer backbone [26, 27, 28], while some other approaches design query-based transformer decoder for semantic segmentation [29, 30].
@@ IV-A Datasets and Evaluation Metrics
NYUv2. The NYUv2 dataset [13] is a typical depth estimation dataset collected from a variety of indoor scenes. The images and corresponding depth maps are captured by the pairs of RGB and Depth cameras. There are 24 images for training, and 654 images for testing. We adopt three commonly used indicators for depth estimation evaluation, including absolute relative error (REL), root mean squared error (RMSE), and average error between predicted depth and ground-truth depth . We further present threshold accuracy which represents =% of pixels satisfying for .
@@ IV-B Experiment Setups
Depth estimation. We adopt the similar decoder as in [51]. During training, we adopt AdamW as the optimizer, and employ the poly learning rate rate schedule. We randomly crop the images into 480480 pixels, and set the batch size as 24. There are totally 25 epochs The initial learning rate is 5e-4 and the weight decay as 0.1. Because depth estimation dataset does not provide ground-truth labels, we employ BLIP to generate image captions fed to explicit language guidance branch.During inference, we employ the techniques of horizontal flipping and sliding windows as in VPD.
@@
TADP[9] | CVPR’24 | 0.225 | 0.976 | 0.997 | 0.999 | 0.062 | 0.027 |
IEDP (Ours) | - | 0.226 | 0.975 | 0.997 | 0.999 | 0.062 | 0.027 |
Method | Publication | Semantic segmentation (ADE20K) | Depth estimation (NYUv2) | ||
mIoU | Time (s) | RMSE | Time (s) | ||
VPD [8] | ICCV’24 | 53.7 | 0.274 | 0.254 | 0.261 |
TADP [9] | CVPR’24 | 54.8 | 1.255 | 0.225 | 1.236 |
IEDP (Ours) | - | 55.9 | 0.296 | 0.226 | 0.280 |
@@ IV-C State-of-the-art Comparison
Depth estimation. We compare our proposed method with some state-of-the-art methods on NYUv2 dataset in Table II. These state-of-the-art methods are also divided into classical approaches and diffusion-based approaches. Compared to diffusion-based approaches, our IEDP achieves promising performance. For example, DDP [44] has the RMSE score of 0.329, VPD [8] has the RMSE score of 0.254, and our IEDP has the RMSE score of 0.226. Therefore, compared to DDP and VPD, our IEDP has the relative improvements of 31.3% and 11.0% in terms of RMSE. Compared to TADP [9], our proposed IEDP has a comparable performance, but does not requires the heavy model BLIP-2 for generating text prompts and explicit text encoder.
Inference time. We further compare our proposed method with diffusion-based methods in terms of both performance and inference time in Table III. We report the inference time of these methods on single NVIDIA A6000 GPU. Compared to VPD [8] and TADP [9], our IEDP has a better trade-off between performance and inference time. For instance, our IEDP outperforms TADP by 1.1% while running at 4.2 times faster on semantic segmentation.
@@
IV-D Ablation Study
Effectiveness of two proposed modules. We first perform the ablation study to show the effectiveness of implicit and explicit prompt modules. Table IV presents the results of progressively integrating these two modules into the baseline. The baseline adopts the unaligned text prompts as in VPD [8], which has the mIoU score of 48.1%. When replacing the unaligned text prompts with our implicit prompts, it has the mIoU score of 49.8%, outperforming the baseline by 1.7%. When further integrating our explicit prompt module, it has the mIoU score of 52.5%, which outperforms the baseline by 4.4% totally.
Different designs of the adapter. Table V shows the results of different adapter designs in implicit branch, where the adapter aims to convert the visual features of CLIP image encoder to implicit text embeddings for denoising UNet. We give two different designs, including using an MLP layer and using learnable queries to convert visual features. We observe that using learnable queries has the better performance.
Number of learnable queries. Table VI shows the impact of different numbers of learnable queries in implicit branch. We observe that there exists a certain degree of performance degradation when using small number of queries, but the decline is relatively limited. Moreover, compared to the baseline, using different queries all have the improvements. We employ the number of queries as 256 as our default setting.
@@ IV-D Ablation Study
Impact of position embeddings. Table VII shows the impact of using position embeddings in both implicit and explicit branch. When not using the position embeddings, it has 0.2% performance drop on mIoU. Therefore, we add position embeddings in both two branches.
Ground-truth labels vs. BLIP-based captions. In explicit language guidance branch, instead of using ground-truth labels, we can also employ BLIP[59] to automatically generate image captions as text prompts like TADP. Table VIII compares ground-truth labels and BLIP-based captions. We observe that it has better performance using ground-truth labels that are less noisy.
@@ V Conclusion
In this paper, we propose an implicit and explicit language guidance framework for diffusion-based visual perception, named IEDP. Our IEDP introduces an implicit prompt module and an explicit prompt module into text-to-image diffusion model. In implicit prompt module, we employ a frozen CLIP image encoder to directly generate implicit text embeddings and fed these embeddings to diffusion model to condition feature extraction. In explicit prompt module, we utilize the ground-truth labels of training images as explicit text prompts and employ CLIP text encoder to generate text embeddings for diffusion model. The implicit prompt module and explicit prompt module jointly train the diffusion model, and only implicit prompt module is used during inference. Experiments are performed on semantic segmentation and depth estimation tasks to show the efficacy of proposed method. Our proposed method achieves promising performance without using manually unaligned text prompts or automatic text prompts with heavy image caption model.